### v2

@@ 4.2 Error Analysis
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{1}{\left(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{2-\beta/\alpha}}$
$\leq\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\mathbb{Z}^{d}\\ m\leq\theta_{\alpha,\boldsymbol{\gamma},M}(\boldsymbol{h})<m+1\end{subarray}}\frac{1}{(\theta_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}))^{2-\beta/\alpha}}$
$\leq\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}\frac{1}{m^{2-\beta/\alpha}}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\mathbb{Z}^{d}\\ m\leq\theta_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})<m+1\end{subarray}}1$
$=\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}\frac{1}{m^{2-\beta/\alpha}}\left(A_{\alpha,\boldsymbol{\gamma}}(m+1)-A_{\alpha,\boldsymbol{\gamma}}(m)\right)$
$\leq\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}A_{\alpha,\boldsymbol{\gamma}}(m+1)\left(\frac{1}{m^{2-\beta/\alpha}}-\frac{1}{(m+1)^{2-\beta/\alpha}}\right)$
$=\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}\frac{A_{\alpha,\boldsymbol{\gamma}}(m+1)}{m^{2-\beta/\alpha}(m+1)^{2-\beta/\alpha}}\times(2-\beta/\alpha)\int_{m}^{m+1}x^{1-\beta/\alpha}\,\mathrm{d}x$
$\leq(2-\beta/\alpha)\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}\frac{A_{\alpha,\boldsymbol{\gamma}}(m+1)}{m^{2-\beta/\alpha}(m+1)^{2-\beta/\alpha}}\times(m+1)^{1-\beta/\alpha}$
$\leq(2-\beta/\alpha)2^{2-\beta/\alpha}\sum_{m=\lfloor H^{-1}_{M}\rfloor}^{\infty}\frac{A_{\alpha,\boldsymbol{\gamma}}(m+1)}{(m+1)^{3-\beta/\alpha}}.$
Here we note that the condition $\theta_{\alpha,\boldsymbol{\gamma},M}(\boldsymbol{h})<m+1$ is satisfied if there exists an $\boldsymbol{\ell}^{\prime}$ such that $r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})<m+1$ holds. For any $\lambda\in[1/2,\alpha)$, it follows from the definition of $A_{\alpha,\boldsymbol{\gamma}}$ and the Cauchy–Schwarz inequality that, for any choice $\boldsymbol{\ell}^{\prime}$, we have
$\frac{A_{\alpha,\boldsymbol{\gamma}}(m+1)}{(m+1)^{1/(2\lambda)}}$ $=\sum_{\begin{subarray}{c}\boldsymbol{h}\in\mathbb{Z}^{d}\\ \theta_{\alpha,\boldsymbol{\gamma},M}(\boldsymbol{h})<m+1\end{subarray}}\frac{1}{(m+1)^{1/(2\lambda)}}$
$\leq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\mathbb{Z}^{d}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})<m+1\end{subarray}}\frac{1}{(m+1)^{1/(2\lambda)}}$
$\leq\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime}))^{1/(2\lambda)}}$
$\leq\left(\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}))^{1/\lambda}}\right)^{1/2}\left(\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h+\ell^{\prime}}))^{1/\lambda}}\right)^{1/2}$
@@
$=\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right).$
This bound, together with the assumptions $1/2<\lambda<\alpha$ and $\beta<\alpha(1-1/(2\lambda))$, leads to
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{1}{\left(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{2-\beta/\alpha}}$
$\leq(2-\beta/\alpha)2^{2-\beta/\alpha}\sum_{m=\lfloor 1/H_{M}\rfloor}^{\infty}\frac{A_{\alpha,\boldsymbol{\gamma}}(m+1)}{(m+1)^{1/(2\lambda)}}\cdot\frac{1}{(m+1)^{3-\beta/\alpha-1/(2\lambda)}}$
$\leq(2-\beta/\alpha)2^{2-\beta/\alpha}\sum_{m=\lfloor 1/H_{M}\rfloor}^{\infty}\frac{1}{(m+1)^{3-\beta/\alpha-1/(2\lambda)}}\times\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
$\leq(2-\beta/\alpha)2^{2-\beta/\alpha}\int_{\lfloor 1/H_{M}\rfloor-1}^{\infty}\frac{1}{(x+1)^{3-\beta/\alpha-1/(2\lambda)}}\,\mathrm{d}x\times\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
$=\frac{(2-\beta/\alpha)2^{2-\beta/\alpha}}{2-\beta/\alpha-1/(2\lambda)}\cdot\frac{1}{(\lfloor 1/H_{M}\rfloor)^{2-\beta/\alpha-1/(2\lambda)}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
@@
$\leq\frac{(2-\beta/\alpha)2^{4-1/(2\lambda)}}{2-\beta/\alpha-1/(2\lambda)}H_{M}^{2-\beta/\alpha-1/(2\lambda)}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right).$
Therefore, we get an upper bound on $B_{M}^{2}$
@@
$B_{M}^{2}$ $\leq\frac{c(2-\beta/\alpha)2^{4-1/(2\lambda)}}{\beta(2-\beta/\alpha-1/(2\lambda))\tau M}H_{M}^{2-\beta/\alpha-1/(2\lambda)}\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
$\leq\frac{c(2-\beta/\alpha)2^{4-1/(2\lambda)}}{\beta(2-\beta/\alpha-1/(2\lambda))\tau M}\left(\frac{2}{(1-\tau)M}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{2}\right)^{2\lambda-\lambda\beta/\alpha-1/2}$
×$\quad\times\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
$\leq\frac{C_{\alpha,\beta,\lambda,\tau}}{M^{2\lambda-\lambda\beta/\alpha+1/2}}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{4\lambda-2\lambda\beta/\alpha}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha}).$
@@
Let $M,d\in\mathbb{N},\alpha>1/2,\boldsymbol{\gamma}\in[0,1]^{\mathbb{N}}$ and $\tau\in(0,1)$ with $M\geq 4$ be given. Assume that (9) holds. For any fixed $\lambda\in(1/2,\alpha)$ and $\beta\in(0,\min(\alpha(1-1/(2\lambda)),1))$, by choosing $T=M^{\lambda-\lambda\beta/(2\alpha)+1/4}$, we have
$e^{\mathrm{rms}\text{-}L_{2}\text{-}\mathrm{app}}_{d,\alpha,\boldsymbol{\gamma}}(A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}(T)})$
$\leq\frac{1}{M^{\lambda/2-\lambda\beta/(4\alpha)+1/8}}\sqrt{1+C_{\alpha,\beta,\lambda,\tau}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{4\lambda-2\lambda\beta/\alpha}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})},$
@@
Let us consider the case where $\lambda\to\alpha-$ and $\beta\to 0+$. The rate of the worst-case RMSE we obtained is of order $M^{-\alpha/2-1/8+\varepsilon}$ for an arbitrarily small $\varepsilon>0$. Moreover, the upper bound shown above is further bounded independently of the dimension $d$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$ as follows: under the assumption $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$, there exists finite $\tilde{d}$ such that
@@ 5 Lower bound
Here, we prove a lower bound on the worst-case root mean squared $L_{2}$-approximation error $e^{\mathrm{rms}\text{-}L_{2}\text{-}\mathrm{app}}_{d,\alpha,\boldsymbol{\gamma}}(A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}(T)})$ of our randomized lattice-based algorithm. First, we present the following lemma, which is essentially due to Byrenheid et al. [3, Lemma 4]. This result is useful for constructing a fooling function $p_{M}\in H_{d,\alpha,\gamma}$ with $\|p_{M}\|_{d,\alpha,\gamma}=1$, which is challenging because our algorithm involves randomly drawing $N$, $\boldsymbol{z}$, and $\boldsymbol{\Delta}$.
@@
$[M_{N,\boldsymbol{z}}^{*}M_{N,\boldsymbol{z}}]_{\ell,\ell^{\prime}}=\sum_{k=0}^{N-1}\exp(2\pi i(\boldsymbol{h}^{(\ell)}-\boldsymbol{h}^{(\ell^{\prime})})\cdot\boldsymbol{x}_{k})=\begin{cases}N&\text{if }\boldsymbol{h}^{(\ell)}-\boldsymbol{h}^{(\ell^{\prime})}=\boldsymbol{0},\\ 0&\text{otherwise},\end{cases}$ (11)
where ,$\ell,\ell^{\prime}\in\{1,2,\ldots,(\lfloor\sqrt{N}\rfloor+1)^{2}\}$. Thus, we have
@@
$\widehat{p_{M}}(\boldsymbol{h}):=\left\{\begin{array}[]{ll}(r_{\alpha,\gamma}^{2}(\boldsymbol{h})|P_{d}(M)|)^{-1/2}&\text{if }\boldsymbol{h}\in P_{d}(M),\\ 0&\text{otherwise}.\\ \end{array}\right.$
@@ 6 Numerical experiments
In the following, for both test functions, Algorithm 3.1 is executed with fixed parameters $\alpha=2$, $\gamma_{1}=\gamma_{2}=1/3$, and $\tau=2/3$ to randomly construct good generating vectors $\boldsymbol{z}$. The remaining parameter, $M$, which determines the range for the number of points, is always set to a power of $2$. Similarly, we use the same parameters $\alpha$, $\gamma_{1}$, $\gamma_{2}$, and $\tau$ to generate the index set $\mathcal{A}_{d}(T)$, where, as motivated by the result in Corollary 4.6, we set $T=M^{\alpha+1/4}=M^{9/4}$. While the choice of parameters may be improved depending on the test functions, it is often unrealistic to assume that the smoothness and weights of the target function are known in advance. The stability of our randomized lattice-based approximation algorithm in terms of these parameters remains an open question for future work.
@@
We now present the results of our numerical experiments. Figure 1 shows the RMSE of our randomized algorithm $A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}}(f)$ plotted against $M$ on a log-log scale. The left and right panels correspond to the test functions $f_{1}$ and $f_{2}$, respectively. Each panel also includes three reference lines indicating theoretical convergence rates: $M^{-\alpha/2}$, which corresponds to the known upper and lower bounds for the worst-case error of deterministic lattice-based algorithms; $M^{-(\alpha/2+1/8)}$, the upper bound obtained in Section 4; and $M^{-(\alpha/2+1/2)}$, the lower bound obtained in Section 5. Here, $\alpha$ is not the parameter used in Algorithm 3.1, but instead reflects the actual smoothness of the test functions, namely $\alpha=1.5$ for $f_{1}$ and $\alpha=2.5$ for $f_{2}$. The thin shaded area around each RMSE curve represents the 95% confidence interval, computed for the squared $L_{2}$-approximation error over $R=1000$ independent trials for each $M$. The narrow width of the interval relative to the RMSE suggests that the reported values are statistically reliable, and the meaningful digits can be considered trustworthy.
As $M$ increases, the convergence behavior improves and appears to approach the asymptotic rate $M^{-(\alpha/2+1/8)}$ shown in Section 4. However, the observed decay of the error stays clearly above the theoretical lower bound $M^{-(\alpha/2+1/2)}$, providing numerical support for the upper bound rather than the lower bound. Closing the gap between the upper and lower bounds remains an open problem for future research.

### v3

@@ 4.2 Error Analysis
$\leq\frac{c}{\beta\tau M}\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}r_{\alpha,\boldsymbol{\gamma}}^{2}(\boldsymbol{h})\,H_{M}^{2-\beta/\alpha}$
$=\frac{c}{\beta\tau M}H_{M}^{2-\beta/\alpha}\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}r_{\alpha,\boldsymbol{\gamma}}^{2}(\boldsymbol{h}),$
[…]
For the sum over $\boldsymbol{h}$ above, we use [21, Lemma 1] to obtain
@@
$\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}r_{\alpha,\boldsymbol{\gamma}}^{2}(\boldsymbol{h})\leq T\,|\mathcal{A}_{d}(T)|\leq T^{1+1/(2\lambda)}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right),$
which holds for any $1/2<\lambda<\alpha$. Therefore, we get an upper bound on $B_{M}^{2}$ as
@@
$B_{M}^{2}$ $\leq T^{1+1/(2\lambda)}\frac{c}{\beta\tau M}H_{M}^{2-\beta/\alpha}\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
@@
$\leq C_{\alpha,\beta,\lambda,\tau}\frac{T^{1+1/(2\lambda)}}{M^{2\lambda-\lambda\beta/\alpha+1}}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{4\lambda-2\lambda\beta/\alpha}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha}).$
@@
Let $M,d\in\mathbb{N},\alpha>1/2,\boldsymbol{\gamma}\in[0,1]^{\mathbb{N}}$ and $\tau\in(0,1)$ with $M\geq 4$ be given. Assume that (9) holds. For any fixed $\lambda\in[1/2,\alpha)$ and $\beta\in(0,1]$, by choosing $T=M^{(2\lambda-\lambda\beta/\alpha+1)/(2+1/(2\lambda))}$, we have
[…]
$\leq\frac{1}{M^{\lambda(2\lambda-\lambda\beta/\alpha+1)/(4\lambda+1)}}\sqrt{1+C_{\alpha,\beta,\lambda,\tau}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{4\lambda-2\lambda\beta/\alpha}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})},$
@@
Let us consider the case where $\lambda\to\alpha$- and $\beta\to 0+$. The rate of the worst-case RMSE we obtained is of order $M^{-\alpha(2\alpha+1)/(4\alpha+1)+\varepsilon}$ for an arbitrarily small $\varepsilon>0$, which is no worse than $M^{-\alpha/2-1/12+\varepsilon}$ for any $\alpha>1/2$. Moreover, the upper bound shown above is further bounded independently of the dimension $d$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$ as follows: under the assumption $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$, there exists finite $\tilde{d}$ such that
@@ 5 Lower bound
[…]
@@
$M_{N,\boldsymbol{z}}^{*}M_{N,\boldsymbol{z}}]_{\ell,\ell^{\prime}}=\sum_{k=0}^{N-1}\exp(2\pi i(\boldsymbol{h}^{(\ell)}-\boldsymbol{h}^{(\ell^{\prime})})\cdot\boldsymbol{x}_{k})=\begin{cases}N&\text{if }\boldsymbol{h}^{(\ell)}-\boldsymbol{h}^{(\ell^{\prime})}=\boldsymbol{0},\\ 0&\text{otherwise},\end{cases}$ (11)
where $\ell,\ell^{\prime}\in\{1,2,\ldots,(\lfloor\sqrt{N}\rfloor+1)^{2}\}$. Thus, we have
@@
[…]
@@ 6 Numerical experiments
In the following, for both test functions, Algorithm 3.1 is executed with fixed parameters $\alpha=2$, $\gamma_{1}=\gamma_{2}=1/3$, and $\tau=2/3$ to randomly construct good generating vectors $\boldsymbol{z}$. The remaining parameter, $M$, which determines the range for the number of points, is always set to a power of $2$. Similarly, we use the same parameters $\alpha$, $\gamma_{1}$, $\gamma_{2}$, and $\tau$ to generate the index set $\mathcal{A}_{d}(T)$, where, as motivated by the result in Corollary 4.6, we set $T=M^{2\alpha(2\alpha+1)/(4\alpha+1)}=M^{20/9}$. While the choice of parameters may be improved depending on the test functions, it is often unrealistic to assume that the smoothness and weights of the target function are known in advance. The stability of our randomized lattice-based approximation algorithm in terms of these parameters remains an open question for future work.
@@
We now present the results of our numerical experiments. Figure 1 shows the RMSE of our randomized algorithm $A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}}(f)$ plotted against $M$ on a log-log scale. The left and right panels correspond to the test functions $f_{1}$ and $f_{2}$, respectively. Each panel also includes three reference lines indicating theoretical convergence rates: $M^{-\alpha/2}$, which corresponds to the known upper and lower bounds for the worst-case error of deterministic lattice-based algorithms; $M^{-\alpha(2\alpha+1)/(4\alpha+1)}$, the upper bound obtained in Section 4; and $M^{-(\alpha/2+1/2)}$, the lower bound obtained in Section 5. Here, $\alpha$ is not the parameter used in Algorithm 3.1, but instead reflects the actual smoothness of the test functions, namely $\alpha=1.5$ for $f_{1}$ and $\alpha=2.5$ for $f_{2}$. The thin shaded area around each RMSE curve represents the 95% confidence interval, computed for the squared $L_{2}$-approximation error over $R=1000$ independent trials for each $M$. The narrow width of the interval relative to the RMSE suggests that the reported values are statistically reliable, and the meaningful digits can be considered trustworthy.
As $M$ increases, the convergence behavior improves and appears to approach the asymptotic rate $M^{-\alpha(2\alpha+1)/(4\alpha+1)}$ shown in Section 4. However, the observed decay of the error stays clearly above the theoretical lower bound $M^{-(\alpha/2+1/2)}$, providing numerical support for the upper bound rather than the lower bound. Closing the gap between the upper and lower bounds remains an open problem for future research.